--> Gemini North Mirror Receives New Silver Coating During Major Shutdown | Gemini Observatory

Change page style: 

Gemini North Mirror Receives New Silver Coating During Major Shutdown

The project list for a shutdown of this magnitude is extremely large.

Listed below are some of the items that were worked during the 2008 Gemini North shutdown  in addition to the mirror coating:


  • A&G corrective maintenance
  • P2 detector problems
  • A&G WSOD - network cable
  • Looked into noise issues with WFSs
  • Jammed cable in AO PA.
  • Installed new termination in cable chains. Not needed for SF RT.
  • Complete mechanical, optical, and electrical PM done on all 4 modules of the A&G
  • SCS worked on Syncrobus problem. Replaced dead fan, and reflective memory in CEM to fix problem.
  • Mechanical systems group played a major role in successfully moving the mirror around.
  • Serviced the Mirror cell cart replacing o-ring seal.
  • Installed invar blocks to mirror to track precisely the position
  • Installed and commissioned pilgrim bolt permanent hydraulic installation.
  • Performed M1 vs M2 alignment measurements.
  • Resealed Mirror covers
  • Commissioned new fans for the extraction of harmful vapors.
  • Optical Maintenance of GCAL and GMOS.
  • Confirmed SF coating is in good shape.
  • LSE work - time stabilization board, change mounting screws and coolant lines, cleaned optics and rod assemblies, retrofit 3 controller modules on laser, cut and splice TEC power supply on/off signal pair. Better temperature stability.
  • Additional UPS installed and tested.
  • Upgrade to summit network storage capabilities
  • Upgrade to summit external network connection
Instrumentation work:
  • GMOS
    • Cleaned
    • Slit mask mechanism service
    • Omega Temp controller for CCD
    • Additional fiber installed for redundancy(all instruments)
    • Troubleshoot cassette installed fault. Problem found in sub circuit assembly.
    • Investigated loss of index due to GMOS interlock system. Replaced suspect chip
    • Vacuum gauge communications converter.
    • Clean up cabling
    • Science group tour.
  • NIRI
    • Cryo cooler fault before shutdown was traced to a problem with the controller and not the cold heads. Repaired.
    • Cleaned up TE cabling
    • Found two loose connections on GNAAC controller. Continued troubleshooting of noise issues.
    • Filter exchange on all three wheels.
    • Replace bad fan in heat exchanger
    • Improve vacuum gauge Communications
  • NIFS
    • Check and install new power supply to the SDSU controller. Extensive testing will be performed on OI before it is hooked up to the science array.
    • Investigate motion of window cover.
    • Move boot bar to accommodate new power supply
  • Altair
    • Assist in synergy board upgrade by burning new EPROM's. Ordered new boards.
    • Dry lubricate mounting bolts


Gemini Observatory Participants